The following thermochemical equation is for the reaction of hydrogen sulfide(g) with oxygen(g) to form water(g) and sulfur dioxide(g). 2H2S(g) + 3O2(g)2H2O(g) + 2SO2(g) H = -1.04×103 kJ When 13.5 grams of hydrogen sulfide(g) react with excess oxygen(g), kJ of energy are .
Answer #1

Molar mass of H2S,

MM = 2*MM(H) + 1*MM(S)

= 2*1.008 + 1*32.07

= 34.086 g/mol

mass(H2S)= 13.5 g

number of mol of H2S,

n = mass of H2S/molar mass of H2S

=(13.5 g)/(34.086 g/mol)

= 0.3961 mol

from reaction,

when 2 mol CO2 reacts, energy = -1040 KJ

So,

for 0.3961 mol, energy = -1040* 0.3961/2 = 206 KJ

Answer: 206 KJ